# Fixed-Point RNNs: Interpolating from Diagonal to Dense

### v2

@@ 3.2 The Fixed-Point Iteration
Solving fixed-point equations such as $h=f_{\theta}(x,h)$, is perhaps one of the most well-studied problems in mathematics (Granas et al., 2003). In the context of deep learning, the literature on Neural ODEs (Chen et al., 2018) and Deep Equilibrium Models (Bai et al., 2019; Ghaoui et al., 2021) investigates fixed-point methods for implicit parametrizations of neural networks. A straightforward, yet effective method computes the forward pass by simply rolling out the fixed-point iteration. In the context of solving $h^{*}=f_{\theta}(x,h^{*})$, this corresponds to introducing an iteration in depth $h^{\ell}=f_{\theta}(x,h^{\ell-1})$. Denoting $\ell$ as the current iteration in depth (i.e., over the layer dimension), and $t$ as the current iteration in time (i.e., over the sequence dimension), the iteration starts at $h_{t}^{0}=0$ and proceeds with
$h_{t}^{\ell}=\mathbf{\Lambda}_{t}h_{t-1}^{{\ell}}+\mathbf{Q}_{t}\mathbf{B}_{t}x_{t}+(\mathbf{I}-\mathbf{Q}_{t})h_{t}^{\ell-1}.$ (6)
Intuitively, this iteration mixes information with interleaved channel mixing (with $\mathbf{Q}_{t}$) and sequence mixing (with $\mathbf{\Lambda}_{t}$) until convergence towards the hidden states of an implicit dense RNN $F_{\theta}$ (cf. 9(a)).
The difficulty with such an iteration in depth and time is that the recurrent dynamics could explode without proper stabilization. While the recurrence in time can be stabilized with RNN techniques  (Zucchet & Orvieto, 2024) such as an input gate $\mathbf{I}-\mathbf{\Lambda}_{t}$, the recurrence in depth, however, could still diverge if $f_{\theta}(x,h)$ does not have an attracting fixed-point (Granas et al., 2003). In order to design a diagonal linear RNN $f_{\theta}(x,h)$ which is guaranteed to have an attracting fixed-point, we make use of Banach (1922)’s theorem. In our context, the theorem states that $f_{\theta}(x,h)$ converges to a fixed-point from any initialization $h^{0}$ if it has a Lipschitz constant $<1$ in $h$. For a fixed-point RNNs with input gate $\mathbf{I}-\mathbf{\Lambda}$, we present the following theorem:
@@
Let $f_{\theta}(x,h)$ be the diagonal linear RNN with input-independent $\mathbf{\Lambda}$ and $\mathbf{Q}$ $f_{\theta}:(x,h)\mapsto h^{\prime},$ $h^{\prime}_{t}=\mathbf{\Lambda}h^{\prime}_{t-1}+\left(\mathbf{I}-\mathbf{\Lambda}\right)\left(\mathbf{Q}\mathbf{B}_{t}x_{t}+\left(\mathbf{I}-\mathbf{Q}\right)h_{t}\right).$ (7) If $||\mathbf{\Lambda}||_{2}<1$ and $||\mathbf{I}-\mathbf{Q}||_{2}<1$, then $f_{\theta}(x,h)$ has a Lipschitz constant $<1$ in $h$. Proof in App. B.1.
Intuitively, Thm. 3.1 states two conditions for stable parametrization of an implicitly dense RNN $F_{\theta}$: (1) the recurrence in time needs to be coupled with input normalization and contractive (i.e. $\|\mathbf{\Lambda}\|_{2}<1$). (2) The recurrence in depth acting on $h$, i.e. $(\mathbf{I}-\mathbf{Q}_{t})$, needs to be contractive. Together, this guarantees that all sequences $h^{\ell}$ up to $h^{*}$ throughout the fixed-point iteration do not explode without any explicit assumptions on the spectral radius on $\mathbf{A}$ (Arjovsky et al., 2016).
@@ 3.3 Parametrization of $\mathbf{Q}_{t}$ and $\mathbf{\Lambda}_{t}$
To satisfy the assumptions required for expressivity in (Cirone et al., 2024b), the implicit transition matrix $\mathbf{A}_{t}$ and therefore $\mathbf{\Lambda}_{t}$ and $\mathbf{Q}_{t}$ need to be input-controlled (i.e. selective), which could be realized through a linear mapping of the input, i.e. $\mathbf{Q}_{t}=\mathcal{M}(x_{t}):=\text{reshape}(\mathbf{W_{Q}}x_{t})$. However, this presents two challenges: how can stability be guaranteed (c.f. Thm. 3.1) and excessive computational cost due to the $O(d^{3})$ parameters of $\mathbf{W_{Q}}$ be avoided? A straight-forward solution lies in structured matrix representations for both the diagonal transition matrix $\mathbf{\Lambda}_{t}$ and the channel mixer $\mathbf{Q}_{t}$.
Inspired by Helfrich et al. (2018), we aim for $\mathbf{Q}_{t}$ to be approximately norm-preserving and $\mathbf{\Lambda}_{t}$ to control the eigenvalue scale using a parametrization akin to Mamba or Griffin (Gu & Dao, 2023; De et al., 2024) and normalization $(\mathbf{I}-\mathbf{\Lambda}_{t})$. For the channel mixers $\mathbf{Q}_{t}$, we consider the structures:
  • Diagonal Plus Low Rank (DPLR): $\mathbf{Q}_{t}=\mathcal{M}(x_{t}):=\left(\mathbf{I}-\sum_{i=1}^{r}\alpha_{it}\cdot\mathbf{\bar{u}}_{it}\mathbf{\bar{u}}_{it}^{\top}\right)$, for rank $r$.
  • Householder Reflections (H): $\mathbf{Q}_{t}=\mathcal{M}(x_{t}):=\prod_{i=1}^{r}\left(\mathbf{I}-\alpha_{it}\cdot\mathbf{\bar{u}}_{it}\mathbf{\bar{u}}_{it}^{\top}\right)$, for $r$ reflections.
  • Kronecker (K): $\mathbf{Q}_{t}=\mathcal{M}(x_{t}):=\mathbf{I}-(\mathbf{\bar{K}}_{t}^{1}\otimes\mathbf{\bar{K}}_{t}^{2})$, where $\otimes$ denotes the Kronecker product.
This allows to reduce the size of the input-dependent parameters $\alpha_{it}$, $\mathbf{\bar{u}}_{it}$, and $\mathbf{\bar{K}}_{t}^{i}$ to $O(d)$, and consequently reduce the size of the linear map $\mathbf{W_{Q}}$ to $O(d^{2}r)$ and $O(d^{2})$. In order to guarantee stability, the condition $||\mathbf{I}-\mathbf{Q}||_{2}<1$ can be enforced by scaling $\alpha_{it}$, $\mathbf{\bar{u}}_{it}$, and $\mathbf{\bar{K}}_{t}^{i}$ appropriately. For more details about the channel mixer variants and how to scale them to negative eigenvalues, please refer to App. C. In Fig. 3(b), we compare different channel mixer variants and observe that the Kronecker structure seems to be most appropriate the state-tracking task $A_{5}$.
@@ 3.4 Algorithmic Implications
This insight suggests an opportunity to introduce a non-linear computation for every sequential step, like in classical RNNs. Concretely, we investigate channel mixers $\mathcal{M}(x_{t}+\smash{h_{t-1}^{\ell-1}})$ which are a function of both the input $x_{t}$ and the hidden state $\smash{h_{t-1}^{\ell-1}}$ from the previous iteration (in both time and depth) without degrading parallelizability. In Fig. 3(b), we compare channel mixers with and without hidden state dependence and observe that this indeed improves sequence length generalization.
@@
$h_{t}^{\ell}=\lambda_{t}^{\ell}\odot h_{t-1}^{{\ell}}+(\mathbf{1}-\lambda_{t}^{\ell})\odot\big{(}\mathbf{Q}_{t}^{\ell}\mathbf{B}_{t}^{\ell}x_{t}+(\mathbf{I}-\mathbf{Q}_{t}^{\ell})h_{t}^{\ell-1}\big{)},$ (8)
where we use $\odot$ to highlight the parallelizability of the element-wise product. We would like to note that due to the normalization $(\mathbf{I}-\mathbf{\Lambda}_{t})$, the corresponding dense RNN $F_{\theta}$ is not explicitly representable anymore. Furthermore, for the time-varying parametrization in Eq. 8, the convergence guarantees may be weaker and solutions $h^{*}$ could be non-unique due to the hidden state dependence. In practice, we iterate until $\frac{||h^{\ell}-h^{\ell-1}||_{\infty}}{||h^{\ell}||_{\infty}}<0.1$ and observe that the conditions of Thm. 3.1 are strong enough to reach convergence as evidenced by Fig. 3(c). Interestingly, the model navigates the parallelism tradeoff (Merrill & Sabharwal, 2023b) and adaptively increases its sequential computation for harder tasks.
@@ 3.5 Optimizing Fixed-Point RNNs
One advantage of converging to a fixed-point as opposed to general layer looping lies in model training. Since the gradient with respect to $h^{0}$ is not needed, implicit differentiation can be used to avoid storing and backpropagating through the computational graph of the fixed-point iteration, as discussed by Liao et al. (2018), Bai et al. (2019), and in App. B.2. In practice, truncated backpropagation of the last $k$ iterations suffices to approximate the gradient through the full iteration $\mathbf{J}_{x}^{*}\approx\mathbf{J}_{x}(h^{\ell^{*}-k})\cdot\ldots\cdot\mathbf{J}_{x}(h^{\ell^{*}})$. For Fixed-Point RNNs we observe that computing the gradient only at the fixed-point ($k=0$), is enough to stabilize training. This means that compared to a single diagonal RNN layer, Fixed-Point RNNs incur no memory overhead and only sequential overhead in the forward pass but not in the backward pass.
We hypothesize that this is possible because $f_{\theta}(x,h)$ is a mostly linear object as opposed to multi-layer implicit models such as (Schöne et al., 2025). Furthermore, we observe that hidden state dependence $\mathcal{M}(x_{t}+h_{t-1}^{\ell-1})$ particularly helps with gradient-based optimization. We credit this to the symmetry between the gradients w.r.t. $x$ and $h$, and formalize this in the following theorem:
@@
Let $f_{\theta}(x,h)$ have Lipschitz constant $<1$ and fixed-point $h^{*}$. If the Jacobians $\frac{\partial f_{\theta}}{\partial x}(x,h)$ and $\frac{\partial f_{\theta}}{\partial h}(x,h)$ are equal, then the gradient $\nabla_{\theta}\mathcal{L}\big{(}f_{\theta}(x,h),y\big{)}$ of the loss $\mathcal{L}(\cdot,y)$ for a target $y$ at the fixed point $h=h^{*}$ will be a descent direction of $\mathcal{L}\big{(}F_{\theta}(x),y\big{)}$. Proof in App. B.3.
@@ 4 Fixed-Point Mamba
In the previous section we introduced the FP-RNN framework on a small RNN with vector hidden state. Now, we extend it to modern matrix state RNNs in Sec. 4.1 and parametrize a dense variant of Mamba (Gu & Dao, 2023) in Sec. 4.2. A detailed description of the architecture is available in App. C.2. We compare the architecture to the baselines Mamba (Gu & Dao, 2023), Mamba-2 (Dao & Gu, 2024), Gated DeltaNet (Yang et al., 2025), and LSTM (Hochreiter & Schmidhuber, 1997) on the copy task introduced by Jelassi et al. (2024) in Sec. 4.3 and state-tracking introduced by Merrill & Sabharwal (2023b) in Sec. 4.4. In order to keep the number of layers at the same order of magnitude, we use two layers (2L) for the diagonal linear RNN baselines and one layer (1L) for FP-Mamba and LSTM. Finally, we discuss the required number of fixed-point iterations in the context of state-tracking and language modeling in Sec. 4.5.
@@ 4.1 Introducing Matrix States
Memory capacity is an important consideration in RNNs. In preliminary experiments, we notice a clear gap between the performance of a Fixed-Point RNNs and Mamba in terms of copying ability. We attribute this difference in performance to Mamba’s state-expansion which endows it with matrix hidden states similar to linear attention, DeltaNet, or mLSTM  (Katharopoulos et al., 2020; Schlag et al., 2021; Beck et al., 2024). In simple terms, these models use an outer product of an input-dependent vector $\mathbf{B}_{t}\in\mathbb{R}^{d_{\text{state}}}$ (i.e. the key) and the input vector $x_{t}\in\mathbb{R}^{d_{\text{inner}}}$ (i.e. the value) as an input to a matrix-valued recurrence with hidden state and transition gate $\mathbf{h}_{t},\boldsymbol{\lambda}_{t}\in\mathbb{R}^{d_{\text{state}}\times d_{\text{inner}}}$. The hidden state is then contracted with another input-dependent vector $\mathbf{C}_{t}\in\mathbb{R}^{d_{\text{state}}}$ (i.e. the query) to get the output $y_{t}^{\top}=\mathbf{C}_{t}^{\top}\mathbf{h}_{t}\in\mathbb{R}^{d_{\text{inner}}}$:
$\mathbf{h}_{t}=\boldsymbol{\lambda}_{t}\odot\mathbf{h}_{t-1}+\mathbf{B}_{t}x_{t}^{\top},$ (9)
This matrix-valued recurrence introduces some challenges to our fixed-point framework. Specifically, in order to mix all the channels over the entirety of the state elements, the mixer $\mathbf{Q}_{t}$ has to be a fourth-order tensor $\in\mathbb{R}^{d_{\text{state}}\times d_{\text{inner}}\times d_{\text{state}}\times d_{\text{inner}}}$ in
$\mathbf{h}_{t}^{\ell}=\boldsymbol{\lambda}_{t}\odot\mathbf{h}_{t-1}^{\ell}+\mathbf{Q}_{t}\otimes\mathbf{B}_{t}x_{t}^{\top}+\left(\mathbf{I}_{d^{4}}-\mathbf{Q}_{t}\right)\otimes\mathbf{h}_{t}^{\ell-1},$ (10)
where, $\otimes$ denotes the fourth-order tensor product and $\mathbf{I}_{d^{4}}$ a fourth-order identity tensor of the same shape as $\mathbf{Q}_{t}$. Certainly computing the fixed-point introduced in Eq. 10 is very challenging both in terms of computation and memory. As we will confirm in Sec. 4.2, one solution is to pass the contracted output $y_{t}$ between fixed-point iterations
$\mathbf{h}_{t}^{\ell}=\boldsymbol{\lambda}_{t}\odot\mathbf{h}_{t-1}^{\ell}+\mathbf{B}_{t}\left(\mathbf{Q}_{t}x_{t}\right)^{\top}+\mathbf{B}_{t}\left(\left(\mathbf{I}-\mathbf{Q}_{t}\right)y_{t}^{\ell-1}\right)^{\top}.$ (11)
This implicitly factorizes the mixer $\mathbf{Q}_{t}$ into separately mixing along dimension $d_{\text{inner}}$ which is used for better expressivity, and dimension $d_{\text{state}}$ which is used for better memory capacity.

### v3

@@ 3.2 The Fixed-Point Iteration
Solving fixed-point equations such as $\mathbf{h}=f_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{h})$, is perhaps one of the most well-studied problems in mathematics (Granas et al., 2003). In the context of deep learning, the literature on Neural ODEs (Chen et al., 2018) and Deep Equilibrium Models (Bai et al., 2019; Ghaoui et al., 2021) investigates fixed-point methods for implicit parametrizations of neural networks. A straightforward, yet effective method computes the forward pass by simply rolling out the fixed-point iteration. In the context of solving $\mathbf{h}^{*}=f_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{h}^{*})$, this corresponds to introducing an iteration in depth $\mathbf{h}^{\ell}=f_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{h}^{\ell-1})$. Denoting $\ell$ as the current iteration in depth (i.e., over the layer dimension), and $t$ as the current iteration in time (i.e., over the sequence dimension), the iteration starts at $\mathbf{h}_{t}^{0}=0$ and proceeds with
$\mathbf{h}_{t}^{\ell}=\mathbf{\Lambda}_{t}\mathbf{h}_{t-1}^{{\ell}}+\mathbf{Q}_{t}\mathbf{B}_{t}\mathbf{x}_{t}+(\mathbf{I}-\mathbf{Q}_{t})\mathbf{h}_{t}^{\ell-1}.$ (6)
Intuitively, this iteration mixes information with interleaved channel mixing (with $\mathbf{Q}_{t}$) and sequence mixing (with $\mathbf{\Lambda}_{t}$) until convergence towards the hidden states of an implicit dense RNN $F_{\boldsymbol{\theta}}$ (cf. 3(a)).
The difficulty with such an iteration in depth and time is that the recurrent dynamics could explode without proper stabilization. While the recurrence in time can be stabilized with RNN techniques  (Zucchet & Orvieto, 2024) such as an input gate $\mathbf{I}-\mathbf{\Lambda}_{t}$, the recurrence in depth, however, could still diverge if $f_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{h})$ does not have an attracting fixed-point (Granas et al., 2003).In order to design a diagonal linear RNN $f_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{h})$ which is guaranteed to have an attracting fixed-point, we make use of Banach (1922)’s theorem. In our context, the theorem states that $f_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{h})$ converges to a fixed-point from any initialization $\mathbf{h}^{0}$ if it has a Lipschitz constant <1<1 in $\mathbf{h}$. For a fixed-point RNNs with input gate $\mathbf{I}-\mathbf{\Lambda}$, we present the following theorem:
@@
Let $f_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{h})$ be the diagonal linear RNN with input-independent $\mathbf{\Lambda}$ and $\mathbf{Q}$ $f_{\boldsymbol{\theta}}:(\mathbf{x},\mathbf{h})\mapsto\mathbf{h}^{\prime},$ $\mathbf{h}^{\prime}_{t}=\mathbf{\Lambda}\mathbf{h}^{\prime}_{t-1}+\left(\mathbf{I}-\mathbf{\Lambda}\right)\left(\mathbf{Q}\mathbf{B}_{t}\mathbf{x}_{t}+\left(\mathbf{I}-\mathbf{Q}\right)\mathbf{h}_{t}\right).$ (7) If $\mathbf{\Lambda}||_{2}<1$ and $\mathbf{I}-\mathbf{Q}||_{2}<1$, then $f_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{h})$ has a Lipschitz constant <1<1 in $\mathbf{h}$. Proof in App. B.1.
Intuitively, Thm. 3.1 states two conditions for stable parametrization of an implicitly dense RNN $F_{\boldsymbol{\theta}}$: (1) the recurrence in time needs to be coupled with input normalization and contractive (i.e. $\|\mathbf{\Lambda}\|_{2}<1$). (2) The recurrence in depth acting on $\mathbf{h}$, i.e. $(\mathbf{I}-\mathbf{Q}_{t})$, needs to be contractive. Together, this guarantees that all sequences $\mathbf{h}^{\ell}$ up to $\mathbf{h}^{*}$ throughout the fixed-point iteration do not explode without any explicit assumptions on the spectral radius on $\mathbf{A}$ (Arjovsky et al., 2016).
@@ 3.3 Parametrization of $\mathbf{Q}_{t}$ and $\mathbf{\Lambda}_{t}$
To satisfy the assumptions required for expressivity in (Cirone et al., 2024b), the implicit transition matrix $\mathbf{A}_{t}$ and therefore $\mathbf{\Lambda}_{t}$ and $\mathbf{Q}_{t}$ need to be input-controlled (i.e. selective), which could be realized through a linear mapping of the input, i.e. $\mathbf{Q}_{t}=\mathcal{M}(\mathbf{x}_{t}):=\text{reshape}(\mathbf{W_{Q}}\mathbf{x}_{t})$. However, this presents two challenges: how can stability be guaranteed (c.f. Thm. 3.1) and excessive computational cost due to the $O(d^{3})$ parameters of $\mathbf{W_{Q}}$ be avoided? A straight-forward solution lies in structured matrix representations for both the diagonal transition matrix $\mathbf{\Lambda}_{t}$ and the channel mixer $\mathbf{Q}_{t}$.
Inspired by Helfrich et al. (2018), we aim for $\mathbf{Q}_{t}$ to be approximately norm-preserving and $\mathbf{\Lambda}_{t}$ to control the eigenvalue scale using a parametrization akin to Mamba or Griffin (Gu & Dao, 2024; De et al., 2024) and normalization $(\mathbf{I}-\mathbf{\Lambda}_{t})$. For the channel mixers $\mathbf{Q}_{t}$, we consider the structures:
  • Diagonal Plus Low Rank (DPLR): $\mathbf{Q}_{t}=\mathcal{M}(\mathbf{x}_{t}):=\left(\mathbf{I}-\sum_{i=1}^{r}\alpha_{it}\cdot\mathbf{\bar{u}}_{it}\mathbf{\bar{u}}_{it}^{\top}\right)$, for rank $r$.
  • Householder Reflections (H): $\mathbf{Q}_{t}=\mathcal{M}(\mathbf{x}_{t}):=\prod_{i=1}^{r}\left(\mathbf{I}-\alpha_{it}\cdot\mathbf{\bar{u}}_{it}\mathbf{\bar{u}}_{it}^{\top}\right)$, for $r$ reflections.
  • Kronecker (K): $\mathbf{Q}_{t}=\mathcal{M}(\mathbf{x}_{t}):=\mathbf{I}-(\mathbf{\bar{K}}_{t}^{1}\otimes\mathbf{\bar{K}}_{t}^{2})$, where $\otimes$ denotes the Kronecker product.
This allows to reduce the size of the input-dependent parameters $\alpha_{it}$, $\mathbf{\bar{u}}_{it}$, and $\mathbf{\bar{K}}_{t}^{i}$ to $O(d)$, and consequently reduce the size of the linear map $\mathbf{W_{Q}}$ to $O(d^{2}r)$ and $O(d^{2})$. In order to guarantee stability, the condition $\mathbf{I}-\mathbf{Q}||_{2}<1$ can be enforced by scaling $\alpha_{it}$, $\mathbf{\bar{u}}_{it}$, and $\mathbf{\bar{K}}_{t}^{i}$ appropriately. For more details about the channel mixer variants, please refer to App. C. Fig. 3(b), we compare different channel mixer variants and observe that the Kronecker structure seems to be most appropriate the state-tracking task $A_{5}$.
@@ 3.4 Algorithmic Implications
This insight suggests an opportunity to introduce a non-linear computation for every sequential step, like in classical RNNs. Concretely, we investigate channel mixers $\mathcal{M}(\mathbf{x}_{t}+\smash{\mathbf{h}_{t-1}^{\ell-1}})$ which are a function of both the input $\mathbf{x}_{t}$ and the hidden state $\smash{\mathbf{h}_{t-1}^{\ell-1}}$ from the previous iteration (in both time and depth) without degrading parallelizability. In Fig. 3(b), we compare channel mixers with and without hidden state dependence and observe that this indeed improves sequence length generalization.
@@
$\mathbf{h}_{t}^{\ell}=\boldsymbol{\lambda}_{t}^{\ell}\odot\mathbf{h}_{t-1}^{{\ell}}+(\mathbf{1}-\boldsymbol{\lambda}_{t}^{\ell})\odot\big(\mathbf{Q}_{t}^{\ell}\mathbf{B}_{t}^{\ell}\mathbf{x}_{t}+(\mathbf{I}-\mathbf{Q}_{t}^{\ell})\mathbf{h}_{t}^{\ell-1}\big),$ (8)
[…]
@@ 3.5 Optimizing Fixed-Point RNNs
One advantage of converging to a fixed-point as opposed to general layer looping lies in model training. Since the gradient with respect to $\mathbf{h}^{0}$ is not needed, implicit differentiation can be used to avoid storing and backpropagating through the computational graph of the fixed-point iteration, as discussed by Liao et al. (2018), Bai et al. (2019), and in App. B.3. In practice, truncated backpropagation of the last $k$ iterations suffices to approximate the gradient through the full iteration $\mathbf{J}_{\mathbf{x}}^{*}\approx\mathbf{J}_{\mathbf{x}}(\mathbf{h}^{\ell^{*}-k})\cdot\ldots\cdot\mathbf{J}_{\mathbf{x}}(\mathbf{h}^{\ell^{*}})$. For Fixed-Point RNNs we observe that computing the gradient only at the fixed-point ($k=0$), is enough to stabilize training. This means that compared to a single diagonal RNN layer, Fixed-Point RNNs incur no memory overhead and only sequential overhead in the forward pass but not in the backward pass.
We hypothesize that this is possible because $f_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{h})$ is a mostly linear object as opposed to multi-layer implicit models such as (Schöne et al., 2025). Furthermore, we observe that hidden state dependence $\mathcal{M}(\mathbf{x}_{t}+\mathbf{h}_{t-1}^{\ell-1})$ particularly helps with gradient-based optimization. We credit this to the symmetry between the gradients w.r.t. $\mathbf{x}$ and $\mathbf{h}$, and formalize this in the following theorem:
@@
Let $f_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{h})$ have Lipschitz constant <1<1 and fixed-point $\mathbf{h}^{*}$. If the Jacobians $\frac{\partial f_{\boldsymbol{\theta}}}{\partial\mathbf{x}}(\mathbf{x},\mathbf{h})$ and $\frac{\partial f_{\boldsymbol{\theta}}}{\partial\mathbf{h}}(\mathbf{x},\mathbf{h})$ are equal, then the gradient $\nabla_{\boldsymbol{\theta}}\mathcal{L}\big(f_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{h}),\mathbf{y}\big)$ of the loss $\mathcal{L}(\cdot,\mathbf{y})$ for a target $\mathbf{y}$ at the fixed point $\mathbf{h}=\mathbf{h}^{*}$ is a descent direction of $\mathcal{L}\big(F_{\boldsymbol{\theta}}(\mathbf{x}),\mathbf{y}\big)$. Proof in App. B.4.
@@ 4 Fixed-Point Mamba
In the previous section we introduced the FP-RNN framework on a small RNN with vector hidden state. Now, we extend it to modern matrix state RNNs in Sec. 4.1 and parametrize a dense variant of Mamba (Gu & Dao, 2024) in Sec. 4.2. A detailed description of the architecture is available in App. C.2. We compare the architecture to the baselines Mamba (Gu & Dao, 2024), Mamba-2 (Dao & Gu, 2024), Gated DeltaNet (Yang et al., 2025), and LSTM (Hochreiter & Schmidhuber, 1997) on the copy task introduced by Jelassi et al. (2024) in Sec. 4.3 and state-tracking introduced by Merrill & Sabharwal (2023) in Sec. 4.4. In order to keep the number of layers at the same order of magnitude, we use two layers for the diagonal linear RNN baselines and one layer for FP-Mamba and LSTM. Finally, we discuss the required number of fixed-point iterations in the context of state-tracking and language modeling in Sec. 4.5.
@@ 4.1 Introducing Matrix States
Memory capacity is an important consideration in RNNs. In preliminary experiments, we notice a clear gap between the performance of a Fixed-Point RNNs and Mamba in terms of copying ability. We attribute this difference in performance to Mamba’s state-expansion which endows it with matrix hidden states similar to linear attention, DeltaNet, or mLSTM  (Katharopoulos et al., 2020; Schlag et al., 2021a; Beck et al., 2024). In simple terms, these models use an outer product of an input-dependent vector $\mathbf{b}_{t}\in\mathbb{R}^{d_{\text{state}}}$ (i.e. the key) and the input vector $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{inner}}}$ (i.e. the value) as an input to a matrix-valued recurrence with hidden state and transition gate $\mathbf{H}_{t},\boldsymbol{\lambda}_{t}\in\mathbb{R}^{d_{\text{state}}\times d_{\text{inner}}}$. The hidden state is then contracted with another input-dependent vector $\mathbf{c}_{t}\in\mathbb{R}^{d_{\text{state}}}$ (i.e. the query) to get the output $\mathbf{y}_{t}^{\top}=\mathbf{c}_{t}^{\top}\mathbf{H}_{t}\in\mathbb{R}^{d_{\text{inner}}}$:
$\mathbf{H}_{t}=\boldsymbol{\lambda}_{t}\odot\mathbf{H}_{t-1}+\mathbf{b}_{t}\mathbf{x}_{t}^{\top},$ (9)
This matrix-valued recurrence introduces some challenges to our fixed-point framework. Specifically, in order to mix all the channels over the entirety of the state elements, the mixer has to be a fourth-order tensor $\mathcal{Q}_{t}\in\mathbb{R}^{d_{\text{state}}\times d_{\text{inner}}\times d_{\text{state}}\times d_{\text{inner}}}$ in
$\mathbf{H}_{t}^{\ell}=\boldsymbol{\lambda}_{t}\odot\mathbf{H}_{t-1}^{\ell}\mathcal{Q}_{t}\bullet\mathbf{b}_{t}\mathbf{x}_{t}^{\top}+\left(\mathcal{I}-\mathcal{Q}_{t}\right)\bullet\mathbf{H}_{t}^{\ell-1},$ (10)
where $\bullet$ denotes the tensor contraction $\mathrm{einsum}(klij,ij\to kl)$ with fourth-order identity tensor $\mathcal{I}$ of the same shape as $\mathcal{Q}_{t}$. Certainly, computing the fixed-point introduced in Eq. 10 is very challenging both in terms of computation and memory. As we will confirm in Sec. 4.2, one solution is to pass the contracted output $y_{t}$ between fixed-point iterations:
$\mathbf{H}_{t}^{\ell}=\boldsymbol{\lambda}_{t}\odot\mathbf{H}_{t-1}^{\ell}+\mathbf{b}_{t}\left(\mathbf{Q}_{t}\mathbf{x}_{t}\right)^{\top}+\mathbf{b}_{t}\left(\left(\mathbf{I}-\mathbf{Q}_{t}\right)\mathbf{y}_{t}^{\ell-1}\right)^{\top}.$ (11)
This implicitly factorizes the tensor mixer $\mathcal{Q}_{t}$ into separately mixing along dimension $d_{\text{inner}}$ which is used for better expressivity, and dimension $d_{\text{state}}$ which is used for better memory capacity.